solve by the addition-or-subtraction method: 12a-5b=18 and 15a-5b=30
Answer by ewatrrr(24785)

 
Hi
solving by the addition-or-subtraction method:
12a-5b=18 and
15a-5b=30
12a-5b= 18
-15a+5b= -30 |mulitply 2nd EQ by -1 to eliminate b
-3a = -12
a = 4 and b = 6 (b = (15*4-30)/5 = 6
CHECKING our Answer***
12a-5b=18
48-30 = 18